OpenStudy (anonymous):

A flower garden is in the shape of a right triangle. The longest side of the triangle measures 13 m. One of the shorter sides is 7 m longer than the other. Find the length on the shorter side

OpenStudy (anonymous):

Do you know Pythagorean Theorem

OpenStudy (anonymous):

a^2+b^2=c^2

OpenStudy (anonymous):

I know, I just want the answer.

OpenStudy (anonymous):

x^2+(x+7)^2=13^2

OpenStudy (anonymous):

It can be completed with complete the square...

OpenStudy (anonymous):

It's on a complete the square worksheet.

OpenStudy (anonymous):

so, that's probably how I should solve it.

OpenStudy (anonymous):

god bless